# Generative AI for Data Augmentation in Wireless Networks: Analysis, Applications, and Case Study

### v1

@@ IV-D Lesson Learned
Current wireless applications may lack sufficient data to support DL model training to meet performance requirements. Generative data augmentation is transforming wireless networks from the physical, network, and application layers. A unified GenAI model can generate jointly data for multiple layers coherently. This cross-layer or multi-layer data augmentation can improve DL models used in communications and networking.
Figure 3: A case study on evaluating the effectiveness of the proposed GDM-based data augmentation framework for Wi-Fi gesture recognition. Part A shows the experiment design, where figures are modified based on [15]. Part B presents the generative performance of the transformer-based diffusion model. Part C illustrates the impact of augmented data on the model performance for Wi-Fi gesture recognition.
V Case Study: Generative Diffusion Models for Data Augmentation in Wi-Fi Gesture Recognition
In this section, we propose a GDM-based data augmentation framework for Wi-Fi gesture recognition and conduct a case study based on a real dataset to evaluate the effectiveness of the proposed framework.
@@ V-A Background Description
Due to its advantages in ubiquitous deployment and non-invasive sensing, Wi-Fi-based gesture recognition is an emerging and innovative application for constructing privacy-preserving and contact-free wireless sensing systems[8]. The use of DL to develop Wi-Fi-based gesture recognition systems has recently garnered significant attention[5, 8], and is not yet mature. One of the biggest challenges in driving the development of this technology is building comprehensive wireless sensing datasets for model training[5]. However, since Wi-Fi sensing is heavily dependent on the surrounding environment, collecting and labeling Wi-Fi sensing data (e.g., CSI data) is costly and energy-intensive[5]. Besides, due to the fundamental differences in wireless data structures, traditional data augmentation methods may not be applicable to Wi-Fi sensing data[5]. To this end, we propose a general GDM-based data augmentation framework for Wi-Fi gesture recognition, as shown in Fig. 2. Specifically, GDMs are leveraged to increase the diversity of Wi-Fi sensing datasets. Then, the raw datasets with the synthesized CSI data can be used to train Deep Neural Network (DNN) models to evaluate the impact of data augmentation.
V-B Proposed Framework
The proposed framework showcases how GDMs can implement CSI data augmentation and boost the performance of Wi-Fi-based gesture recognition systems, which is composed of three parts:
V-B1 Pre-processing raw CSI data
@@
In Wi-Fi sensing systems, CSI is estimated on multiple subcarriers[5], providing environmental information on how each Wi-Fi subcarrier is affected during transmission. CSI features lay the foundation of wireless sensing, including the amplitude and phase of signals transmitted between Wi-Fi devices. Before CSI data synthesis by GDMs, it is necessary to pre-process the original CSI data based on feature extraction, as shown in the left part of Fig. 2. The processed data consists of CSI features and condition information, where the features are represented as a complex matrix in double precision, and the condition information is used for conditional generation, including the gestures, orientations, and locations of users.
V-B2 Synthesizing high-quality CSI data
@@
Based on the processed CSI data, we use the transformer-based diffusion model to synthesize high-quality CSI data[8], as shown in the middle part of Fig. 2. In the forward diffusion process, time-series noise is progressively added to the input data until following the Gaussian distribution. Then, the diffusion model utilizes the denoising network to iteratively eliminate the noise and restore the original data distribution. The reverse restoration process is designed with a hierarchical architecture, which can fully reveal the time-frequency characteristics of CSI data. Moreover, the denoising network is essentially a complex-valued neural network with an embedded Transformer module to further extract CSI data features, thus generating high-quality and time-series CSI data.
V-B3 Evaluating augmented data
@@
As shown in the right part of Fig. 2, we use DNN models to evaluate the impact of augmented data on the performance of Wi-Fi-based gesture recognition systems. Nowadays, Doppler Frequency Shift (DFS) spectrograms are generally used as inputs to DL models[5]. DFS spectrograms are extracted from CSI data without intensive computation and reflect the movement and activities of the target, which is capable of striking a balance between denoising raw data and preserving signals for learning[5]. Thus, we transform both the raw and generated CSI data to DFS spectrograms through Short-Term Fourier Transform (STFT). Specifically, STFT can truncate the time series of signals using a fixed-length sliding time window and perform a fast Fourier transform on each window[5]. Finally, both the original and generated DFS spectrograms can be used in the training of DNN models in the form of images.
@@ V-C Simulation Experiments
We implement the GDM-based data augmentation framework for Wi-Fi gesture recognition by using PyTorch on NVIDIA GeForce RTX 3080 Laptop GPU and evaluate its effectiveness on a real dataset from Widar 3.0[15], where Widar 3.0 is the largest WiFi sensing public dataset[8, 5]. Our project is available here [^2] https://github.com/JinboWen00/GDA-for-Wi-Fi-Gesture-Recognization.
@@ V-C1 Experiment design
The Widar 3.0 dataset is collected from 17 users. Users stand in 5 different locations within the range of 6 receivers equipped with Intel 5300 network cards, oriented in 5 different directions towards a transmitter, and perform 22 gestures[5, 15]. Due to the non-uniform distribution of gesture data, we select one user with corresponding 6 gestures as the dataset for the evaluation of the proposed framework. The experimental scenario is a hall, and the 6 gestures are pushing and pulling, sweeping, clapping, sliding, horizontally drawing O, and horizontally drawing zigzag, respectively, as shown in the Part A of Fig. 3. In addition, we develop a ResNet [^3] https://www.geeksforgeeks.org/residual-networks-resnet-deep-learning/ model to evaluate the impact of augmented CSI data on Wi-Fi gesture recognition performance, where we utilize the Adam algorithm for gradient optimization and set the network learning rate as $0.0001$.
@@ V-C2 Simulation results
As shown in the Part B of Fig. 3, we present the ground truth DFS spectrogram and the generated DFS spectrogram for 6 gestures, along with their corresponding Fréchet Inception Distance (FID) values and average Structural Similarity Index Measure (SSIM) values. The SSIM and the FID are two common image quality indexes that can evaluate the generative quality of transformer-based diffusion models, where the SSIM is a perceptual method to quantify the similarity between two images and the FID is an effective metric used to access the realism and diversity of generated images[8]. Note that the SSIM value is $[-1,1]$, where the closer the value is to $1$, the better the image quality, and the FID value is $[0,\infty)$, where the smaller the value, the closer the distribution of the generated image and the real image. We can observe that the transformer-based diffusion model can generate high-quality DFS spectrograms close to real data distribution.
In the Part C of Fig. 3, we explore the impact of synthesized data ratio. Firstly, since the ResNet is one of the most commonly used backbone CV models[5], we develop a ResNet model for Wi-Fi gesture recognition without data augmentation, and the proposed ResNet model can effectively recognize Wi-Fi gestures, as shown in the result module (C1). To evaluate the impact of synthesized data on the performance of Wi-Fi gesture recognition systems, we use the generated DFS spectrograms with the original DFS spectrograms to retrain the ResNet model. As can be seen from the result module (C2), the average validation accuracy of the ResNet model first increases and then decreases as the proportion of augmented data increases. The reason is that a small amount of augmented data can enrich the diversity of training datasets and improve the performance of DL models. However, as the augmented data gradually increases, the original data distribution may be replaced by the generated data distribution, potentially resulting in model collapse.
@@ VI-A Generative Data Augmentation for Next-Generation Networks
Next-generation networks are expected to support ultra-reliable low-latency communications, massive machine-type communications, and high-precision localization, enabling intelligent and promising applications, such as autonomous vehicles and immersive extended reality. However, these applications may present unique challenges, such as data hungry[14], which cannot be effectively addressed by using traditional DL models and datasets. Therefore, future research can focus on utilizing generative data augmentation to create synthetic data that can enhance traditional datasets and accurately reflect the diverse conditions of these novel applications, thus enabling better DL model training.
@@ VI-B Multi-modal Generative Data Augmentation for Wireless Networks
Traditional data augmentation often focuses on single data modalities. However, the next generation of wireless networks, especially with the advent of six-generation networks, will need to process multiple types of data simultaneously, such as spatial, temporal, and contextual data. Therefore, the realization of multi-modal generative data augmentation technology is significant and promising for future research in data augmentation. The Mixture of Experts (MoE) model is expected to promote the realization of multi-modal generative data augmentation technology. Specifically, in the MoE architecture, each expert can be specialized to handle a specific modality and generate data within its domain.
@@ VI-C High-quality Data Augmented by Generative AI
The effectiveness of DL-based wireless communications and networking largely depends on the quality of training datasets. Low-quality or unrealistic training data can cause models to perform poorly or behave unpredictably in real wireless networks. As GenAI plays an increasingly pivotal role in augmenting data for wireless networks, ensuring the quality of the generated data becomes critically important. Therefore, how to evaluate and guarantee the quality of data augmented by GenAI models is a critical future direction.
@@ VII Conclusion
In this article, we have explored the potential and effectiveness of GenAI-driven data augmentation in wireless networks. We have discussed the limitations of traditional data augmentation methods in wireless networks and introduced generative data augmentation. Then, we have explored GenAI-driven data augmentation for wireless applications from the physical, network, and application layers. Subsequently, we have proposed a general generative diffusion model-based data augmentation framework for Wi-Fi gesture recognition and developed ResNet models to evaluate the impact of augmented CSI data on Wi-Fi gesture recognition performance. Simulation results demonstrate the effectiveness of the proposed framework through case studies based on a real dataset from Widar 3.0. Finally, we have investigated potential open research directions for generative data augmentation.

### v2

@@ IV-D Lesson Learned
Current wireless applications may lack sufficient data to support DL model training to meet performance requirements. Generative data augmentation is transforming wireless networks from the physical, network, and application layers. Unified GenAI models can generate data jointly for multiple layers coherently. This cross-layer or multi-layer data augmentation improves DL models used in communications and networking.
Figure 3: A case study on evaluating the effectiveness of the proposed generative data augmentation framework for Wi-Fi gesture recognition. Part A shows the experiment design, where figures are modified based on [15]. Part B presents the generative performance of the transformer-based diffusion model, which is the SOTA model for RF data augmentation compared with traditional GenAI models such as GAN and VAE models. Part C illustrates the performance of the generative data augmentation approach for Wi-Fi gesture recognition.
V Case Study: Transformer-based Diffusion Models for Data Augmentation in Wi-Fi Gesture Recognition
In this section, we introduce the generative data augmentation framework for Wi-Fi gesture recognition and conduct a case study using the Widar 3.0 dataset to evaluate the effectiveness of the proposed framework.
@@ V-A Background Description
Wi-Fi gesture recognition is an emerging and innovative application for constructing privacy-preserving and contact-free wireless sensing systems [8]. The application of DL to Wi-Fi gesture recognition has recently garnered significant attention [5, 8], but the field remains relatively immature. One of the biggest challenges in driving the development of this technology is building comprehensive wireless sensing datasets for model training [5]. However, since Wi-Fi sensing is heavily dependent on the surrounding environment, collecting and labeling Wi-Fi sensing data (e.g., CSI data) is costly and energy-intensive [5]. Moreover, due to the fundamental differences in wireless data structures, traditional data augmentation methods may not be applicable to Wi-Fi sensing data [5]. Therefore, we propose a general generative data augmentation framework for Wi-Fi gesture recognition, as shown in Fig. 2. Specifically, we leverage transformer-based diffusion models to enhance the diversity of Wi-Fi sensing datasets. The original datasets, augmented with the synthesized data, are then used to train Deep Neural Network (DNN) models to evaluate the impact of data augmentation.
V-B Framework Design
We present the design details and practical deployment of the generative data augmentation framework for Wi-Fi gesture recognition, consisting of three steps:
@@ V-B Framework Design
Step 1. Pre-processing raw CSI data: In Wi-Fi sensing systems, CSI is estimated on multiple subcarriers [5], providing environmental information on how each Wi-Fi subcarrier is affected during transmission, which can capture the fine-grained changes in the channel due to human motion. CSI features lay the foundation of wireless sensing, including the amplitude and phase of signals transmitted between Wi-Fi devices. Before CSI data synthesis, we first pre-process the original CSI data based on feature extraction, as shown in Part 1 of Fig. 2. The processed data consists of CSI features and conditional information, where the features are represented as a complex matrix in double precision, and the conditional information is used for the denoising process, including the gestures, orientations, and locations of users.
@@
Step 2. Synthesizing high-quality CSI data: Based on the processed CSI data, we leverage the pre-trained transformer-based diffusion model, which is the SOTA model for RF data augmentation, to synthesize high-quality CSI data [8]. The architecture of the transformer-based diffusion model is shown in Part 2 of Fig. 2. In the forward diffusion process, time-series noise is progressively added to the original CSI data until it approximates a Gaussian distribution. Then, the diffusion model utilizes the denoising network to iteratively eliminate the noise and restore the original data distribution. The reverse restoration process is designed with a hierarchical architecture, which can fully reveal the time-frequency characteristics of CSI data. Moreover, the denoising network is essentially a complex-valued neural network with an embedded Transformer module to further extract CSI data features, thus generating high-quality and time-series CSI data.
@@
Step 3. Evaluating generative data augmentation: As illustrated in Part 3 of Fig. 2, we use DNN models to evaluate the performance of generative data augmentation for Wi-Fi gesture recognition. Nowadays, Doppler Frequency Shift (DFS) spectrograms are generally used as inputs to DL models [5]. DFS spectrograms are extracted from CSI data without intensive computation and reflect the movement and activities of the target, which is capable of striking a balance between denoising raw data and preserving signals for learning [5]. Hence, we transform both the raw and generated CSI data to DFS spectrograms through Short-Term Fourier Transform (STFT). Specifically, STFT can truncate the time series of signals using a fixed-length sliding time window and perform a fast Fourier transform on each window [5]. Finally, both the original and generated DFS spectrograms can be used in the training of DNN models in the form of images.
TABLE II: Performance Comparisons of ResNet-18 and MobileNetV2 in Generative Data Augmentation.
Metric ResNet-18 MobileNetV2
0% 20% 40% 60% 80% 100% 0% 20% 40% 60% 80% 100%
Accuracy $\uparrow$ 0.8974 0.8989 0.9064 0.9054 0.9056 0.9127 0.8642 0.8602 0.8618 0.8615 0.8641 0.8658
Precision $\uparrow$ 0.8976 0.9004 0.9057 0.9051 0.9042 0.9113 0.8700 0.8668 0.8704 0.8660 0.8677 0.8708
Recall $\uparrow$ 0.8974 0.8989 0.9064 0.9054 0.9055 0.9127 0.8640 0.8602 0.8616 0.8615 0.8640 0.8658
F1-Score $\uparrow$ 0.8936 0.8967 0.9034 0.9033 0.9029 0.9103 0.8601 0.8558 0.8560 0.8543 0.8586 0.8618
@@ V-C Simulation Experiments
We implement the generative data augmentation framework for Wi-Fi gesture recognition by using PyTorch on NVIDIA GeForce RTX 3080 Laptop GPU and evaluate its effectiveness on the Widar 3.0 dataset [15], which is the largest publicly available Wi-Fi sensing public dataset [8, 5].
@@ V-C1 Experiment design
The Widar 3.0 dataset [15] is collected from 17 users, performing 22 distinct gestures. Each user stands at 5 different locations within the range of 6 receivers equipped with an Intel 5300 network card, oriented in 5 different directions towards a transmitter. Because of the non-uniform distribution of gesture data across users, we select two subdatasets consisting of one user and 6 representative gestures for the evaluation of the proposed framework. As shown in the Part A of Fig. 3, these gestures include pushing and pulling, sweeping, clapping, sliding, drawing a circle, and drawing a zigzag. The experimental scenarios involve a hall and an office.
In our implementation, we first pre-process the CSI data using MATLAB. The processed CSI data is then used as the conditional input to a pre-trained transformer-based diffusion model, which generates corresponding DFS spectrograms. To evaluate the impact of the proposed generative data augmentation framework on Wi-Fi gesture recognition performance, we employ a ResNet-18 model with pre-trained ImageNet weights as the feature extractor [5]. During training, the model is optimized using the Adam algorithm with a learning rate of $0.0001$, a batch size of $32$, and $20$ training epochs. To ensure reproducibility, we fix the random seed at $42$.
@@ V-C2 Simulation results
As shown in the Part B of Fig. 3, we present both the ground truth DFS spectrogram and the generated DFS spectrogram for 6 gestures, along with their corresponding Fréchet Inception Distance (FID), average Structural Similarity Index Measure (SSIM), and 1-Wasserstein Distance (WD) values to evaluate the generative quality of transformer-based diffusion models. SSIM is a perceptual metric used to quantify the similarity between a generated image and its reference [8]. Its value ranges from 1-1 to $1$, with higher values indicating better generative quality. Both FID and 1-WD are effective metrics used to assess the distributional divergence between generated and real images [8]. Their values are $0,\infty)$, with lower values indicating better fidelity. From Fig. 3, we observe that the transformer-based diffusion model can generate high-quality DFS spectrograms that closely resemble the real data distribution across the 6 gesture classes. The reason is that transformer-based diffusion models are optimized using a likelihood-based objective [8], which avoids model collapse issues commonly seen in GANs, enabling improved training stability. Moreover, with the role of the attention mechanism, transformer-based diffusion models excel at capturing the global structure of data distributions, thereby generating high-quality DFS spectrograms. It is worth noting that the pre-trained transformer-based diffusion model supports inference on either a CPU or a standard desktop-grade GPU, which generates hundreds of data samples within a few minutes, indicating practical computational efficiency in generative data augmentation.
In the Part C of Fig. 3, we evaluate the effectiveness of generative data augmentation. We first compare the impact of the proposed generative data augmentation approach with that of a traditional data augmentation approach (i.e., cropping) on the performance of the ResNet-18 model. As shown in the result module (C1) of Fig. 3, we observe that the generative data augmentation approach positively contributes to improving the average recognition accuracy, whereas the traditional data augmentation approach has a negative impact. Moreover, as the proportion of augmented data increases, the performance gap between the two approaches becomes increasingly pronounced. The reason is that traditional data augmentation approaches may disrupt the frequency information of the original DFS spectrograms, resulting in the generation of low-quality augmented data. In contrast, only high-quality augmented data can effectively enrich the diversity of training datasets and improve the performance of the ResNet-18 model. We then evaluate the impact of the generative data augmentation approach in two distinct environments: a hall and an office. As shown in the result module (C2) of Fig. 3, the generative data augmentation approach has a positive impact on improving the recognition accuracy in both hall and office environments, highlighting the scalability and robustness of the proposed framework. Finally, we conduct a performance comparison between ResNet-18 and MobileNetV2 under the proposed framework. As illustrated in Table II, we observe that the generative data augmentation approach effectively enhances the performance of both models in Wi-Fi gesture recognition. Moreover, ResNet-18 consistently outperforms MobileNetV2 across all four evaluation metrics, including macro precision, macro recall, and macro F1-score. The reason is that the residual architecture of ResNet-18 enables it to extract more discriminative features from DFS spectrogram images, demonstrating the superiority of ResNet-18 models in Wi-Fi gesture recognition.
V-D Lesson Learned
The above simulation results demonstrate the effectiveness of the proposed generative data augmentation framework. The proposed framework also offers inherent privacy-preserving benefits by synthesizing CSI data that retains task-relevant features, avoiding the additional acquisition of raw CSI data that may contain sensitive information.
@@ VI-A Generative Data Augmentation for Next-Generation Networks
Next-generation networks are expected to support ultra-reliable low-latency communications, massive machine-type communications, and high-precision localization, enabling intelligent and promising applications, such as autonomous vehicles and immersive extended reality. However, these applications may present unique challenges, such as being data hungry [14], which cannot be effectively addressed by using traditional DL models and datasets. Therefore, future research can focus on utilizing generative data augmentation to create synthetic data that can enhance traditional datasets and accurately reflect the diverse conditions of these novel applications, thus enabling better DL model training.
@@ VI-B Multi-modal Generative Data Augmentation for Wireless Networks
Traditional data augmentation often focuses on single data modalities. However, the next generation of wireless networks, especially with the advent of sixth-generation networks, will need to process multiple types of data simultaneously, such as spatial, temporal, and contextual data. Therefore, the realization of multi-modal generative data augmentation technology is significant and promising for future research in data augmentation. The Mixture of Experts (MoE) model is expected to promote the realization of multi-modal generative data augmentation technology. Specifically, in the MoE architecture, each expert can be specialized to handle a specific modality and generate data within its domain.
@@ VI-C High-quality Data Augmented by Generative AI
The effectiveness of DL-based wireless communications and networking largely depends on the quality of training datasets. Low-quality or unrealistic training data can cause models to perform poorly or behave unpredictably in real wireless networks. As GenAI plays an increasingly pivotal role in augmenting data for wireless networks, ensuring the quality of the generated data becomes critically important. Therefore, how to consistently guarantee the quality of data augmented by GenAI models is a critical future direction.
@@ VII Conclusion
In this article, we have explored the potential and effectiveness of generative data augmentation in wireless networks. We have discussed the limitations of traditional data augmentation methods in wireless networks and introduced generative data augmentation. Then, we have explored generative data augmentation for wireless applications from the physical, network, and application layers. Subsequently, we have proposed a general generative data augmentation framework for Wi-Fi gesture recognition. Furthermore, we have conducted a case study using the Widar 3.0 dataset, which employed ResNet-18 models to evaluate the impact of generative data augmentation on Wi-Fi gesture recognition performance. Extensive simulation results demonstrate the effectiveness of the proposed framework. Finally, we have investigated potential open research directions for generative data augmentation.